### 05 Mar 2026 v2

# ORCID 0000-0001-6130-9413

Abstract.
In the paper we describe some properties of function
$y=r(x)=\lim_{n\to\infty}\frac{1}{n}\sum^{\infty}_{k=1}\alpha_{k}(x),\text{ where }x=\sum^{\infty}_{k=1}\alpha_{k}(x)4^{-k}$
of 44-adic digits asymptotic mean of fractional part of real number $x$, particularly properties of it’s level sets $S_{\theta}=\left\{x:r(x)=\theta,\>\theta=const,\>0\leqslant\theta\leqslant 3\right\},$ if all 44-adic digits frequencies exist, i.e.
$\nu_{i}(x)=\lim_{n\to\infty}n^{-1}\#\{k:\alpha_{k}(x)=i,i\leqslant n\},\>\>i=0,1,2,3.$
We provided an algorithm of constructing point from the set $S_{\theta}$, and proved continuality and every where density of the set. We found conditions of zero and full Lebesgue measure and estimates of Hausdorff–Besicovitch fractal dimension.
Key words and phrases:
Asymptotic mean of digits, digit frequency of number, 44–adic representation of real numbers, sets of Besicovitch–Eggleston type, Hausdorff–Besicovitch fractal dimension, normal numbers, weakly normal numbers.
1991 Mathematics Subject Classification:
11K55, 26A27, 26A30
Scientific Journal of Drahomanov National Pedagogical University. Series 1. Physical and Mathematical Sciences. – Kyiv: Drahomanov National Pedagogical University, 2013, No. 14, pp. 217–226.
1. Introduction
We study the fractional part of a real number; therefore, we restrict our consideration to numbers from the interval $[0,1]$. Let $s\geqslant 2$ denote a fixed natural number and let $\mathcal{A}_{s}={0,1,\ldots,s-1}$ denote the alphabet of the ss–adic numeral system. It is well known that for any $x\in[0,1]$ there exists a sequence $(\alpha_{n})$, where $\alpha_{n}\in\mathcal{A}_{s}$, such that
$x=\displaystyle\frac{\alpha_{1}}{s}+\displaystyle\frac{\alpha_{2}}{s^{2}}+\cdots+\displaystyle\frac{\alpha_{n}}{s^{n}}+\cdots\equiv\Delta^{s}_{\alpha_{1}\alpha_{2}\ldots\alpha_{n}\ldots}.$ (1)
The symbolic notation $\Delta^{s}_{\alpha_{1}\alpha_{2}\ldots\alpha_{n}\ldots}$ of series (1) is called ss–adic representation of the number.
All irrational numbers and some rational numbers admit a unique ss–adic representation and we call such numbers ss–adic irrational. The remaining numbers (the set of these numbers is countable) admit exactly two ss–adic representations, namely:
$\Delta^{s}_{c_{1}\ldots c_{k-1}c_{k}(0)}=\Delta^{s}_{c_{1}\ldots c_{k-1}[c_{k}-1](s-1)},$
where $(i)$ denotes the period in the ss–adic representation of the number. We call such numbers ss–adic rational. To define the nn–th digit $\alpha_{n}(x)$ of a number $x$ as a function of $x$, we agree to use only the first ss–adic representation, namely, the one that has period $(0)$.
Using the ss–adic representation of numbers it was defined and studied many different mathematical objects with complex local structure and fractal properties. These include sets, functions, distributions of random variables, dynamical systems, space transformations, and others. We can also use the following concept for the same purposes.
We define the asymptotic mean of digits of the number $x=\Delta^{s}_{\alpha_{1}\alpha_{2}\ldots\alpha_{n}\ldots}$ as the value $r(x)$ given by the limit
$\lim\limits_{n\to\infty}\frac{1}{n}\sum\limits^{n}_{i=1}\alpha_{i}(x)\equiv r(x),$
provided that the limit exists.
We introduced the concept of the asymptotic mean of digits and its application to the study of the topological–metric and fractal properties of sets of real numbers in paper [10].
We focus on the topological–metric properties of sets of numbers with a given asymptotic mean of digits, that is, sets of the form
$S_{\theta}\equiv\left\{x:\lim_{n\to\infty}\frac{1}{n}\sum^{n}_{i=1}\alpha_{i}(x)=\theta\geqslant 0\right\},$
where the constant $\theta$ is a predetermined parameter.
The concept of the asymptotic mean of digits of a number is closely related to the concept of digit frequency. In the case of the binary numeral system these concepts coincide. Let us recall this notion.
Let $N_{i}(x,n)$ denote the quantity of digits $i\in\mathcal{A}s$ in the ss–adic representation $\Delta^{s}{\alpha_{1}\alpha_{2}\ldots\alpha_{k}\ldots}$ of a number $x\in[0,1]$ up to and including the nn–th place, that is,
$N_{i}(x,n)=\#\{j:\,\alpha_{j}(x)=i,\,j\leqslant n\}.$
We define the frequency of the digit $i$ in the ss–adic representation of a number $x\in[0,1]$ as the limit (if it exists)
$\nu_{i}(x)=\lim\limits_{n\to\infty}\displaystyle\frac{N_{i}(x,n)}{n}.$
The frequency function $\nu_{i}(x)$ of the digit $i$ in the ss–adic representation of a number $x\in[0,1]$ is well-defined for ss–adic irrational numbers, and for ss–adic rational numbers it is well-defined after we agree to use only the representation with period $(0)$.
The numer $r_{n}(x)\equiv\frac{1}{n}\sum\limits^{n}_{i=1}\alpha_{i}(x)$ is called the relative mean of the digits of the number $x$. Since
$r_{n}(x)=\displaystyle\frac{N_{1}(x,n)}{n}+\displaystyle\frac{2N_{2}(x,n)}{n}+\ldots+\displaystyle\frac{(s-1)N_{s-1}(x,n)}{n},$
we have $0\leqslant r_{n}(x)\leqslant s-1$. It follows that if the frequencies of all digits exist, then the asymptotic mean of digits also exists.
We call a number $x$ normal on the base $s$ (weakly normal) if for each $i\in\mathcal{A}_{s}$ the frequency exists and equals $\nu_{i}(x)=s^{-1}$.
The set of normal numbers in the interval $[0,1]$ has full Lebesgue measure [4].
We define the Besicovitch–Eggleston set $E[\tau_{0},\tau_{1},\ldots,\tau_{s-1}]$ as
$E[\tau_{0},\tau_{1},\ldots,\tau_{s-1}]=\{x:\nu_{i}(x)=\tau_{i},\,\,i=\overline{0,s-1}\}.$
The Hausdorff–Besicovitch fractal dimension $\alpha_{0}(\cdot)$ of the set $E[\tau_{0},\tau_{1},\ldots,\tau_{s-1}]$ can be calculated [7] using the formula
$\alpha_{0}(E[\tau_{0},\tau_{1},\ldots,\tau_{s-1}])=-\frac{\ln\tau_{0}^{\tau_{0}}\tau_{1}^{\tau_{1}}\ldots\tau_{s-1}^{\tau_{s-1}}}{\ln(s-1)}.$
We focus on the case $s=4$, since the case $s=3$ is analyzed in detail in our paper [10]. In fact, for $s>3$ the set $S_{\theta}$ exhibits richer properties.
We consider the set of numbers with a prescribed asymptotic mean of digits in their 44–adic representation, that is, sets of the form
$S_{\theta}\equiv\left\{x:r(x)=\theta\right\},$
where $\theta$ is a predetermined parameter from the interval $[0,3]$.
The set $S_{\theta}$ is the union of three disjoint sets $\Theta_{1}$, $\Theta_{2}$, and $\Theta_{3}$, such that
[…]
We now analyze the properties of the subset $\Theta_{1}$ of the set $S_{\theta}$.
2. The set $\Theta_{1}$ and the Besicovitch–Eggleston sets
Theorem 1.
If $\theta=0$ or $\theta=3$, then $\Theta_{1}$ is an anomalously fractal and everywhere dense set.
Proof.
Let $v^{(n)}_{j}=n^{-1}N_{j}(x,n)$ denote the relative frequency of the digit $j$ in the 44–adic representation of the number $x$, and let $r_{n}(x)=\frac{1}{n}\sum\limits^{n}_{j=1}\alpha_{j}(x)$ denote the relative mean of the digits of $x$. Then the following system of equations holds:
[…]
Let $\theta=0$. If $\lim\limits_{n\to\infty}r_{n}(x)=0$ then for each $i\in{1,2,3}$ the following condition holds: $0\leqslant v^{(i)}_{n}(x)\leqslant v^{(1)}_{n}(x)+2v^{(2)}_{n}(x)+3v^{(3)}_{n}(x)=r_{n}(x)\to 0$ as $n\to\infty$, hence $\nu_{i}(x)=\lim\limits_{n\to\infty}v^{(i)}_{n}(x)=0$ and respectively $\nu_{0}(x)=1$. Therefore, $S_{\theta}=\Theta_{1}=E[1,0,0,0]$. This set is everywhere dense set and its Hausdorff–Besicovitch dimension is equal to
$\alpha_{0}(E[1,0,0,0])=\frac{\ln 1^{1}0^{0}0^{0}0^{0}}{-\ln 4}=0.$
Now let $\theta=3$. If $\lim\limits_{n\to\infty}r_{n}(x)=3$, then multiplying the first equation of system ()(*) by 3 and subtracting the second equation of the system, we obtain: $3v^{(0)}_{n}+2v^{(1)}_{n}+v^{(2)}_{n}=3-r_{n}$. Hence, $0\leqslant v^{(i)}_{n}(x)\leqslant 3v^{(0)}_{n}(x)+2v^{(1)}_{n}(x)+v^{(2)}_{n}(x)=3-r_{n}(x)\to 0$ as $n\to\infty$. Hence, $\nu_{i}(x)=0$ for all $i\in\{0,1,2\}$ and $\nu_{3}(x)=1$. Therefore, $\Theta_{2}=\Theta_{3}=\varnothing$ and $S_{\theta}=\Theta_{1}=E[0,0,0,1]$. This set is everywhere dense, and its Hausdorff–Besicovitch dimension equals 0. ∎
If the 44–adic representation of a number $x$ has frequencies of all digits $\nu_{0}$, $\nu_{1}$, $\nu_{2}$, $\nu_{3}$, then it has an asymptotic mean of digits $r(x)$, given by equality
$r(x)=\nu_{1}(x)+2\nu_{2}(x)+3\nu_{3}(x).$
Thus, the set $\Theta_{1}$ is the union of Besicovitch–Eggleston sets $E[\tau_{0},\tau_{1},\tau_{2},\tau_{3}]$ over all possible probability vectors $(\tau_{0},\tau_{1},\tau_{2},\tau_{3})$ satisfying $\tau_{1}+2\tau_{2}+3\tau_{3}=\theta$, that is,
$\Theta_{1}=\bigcup E[\tau_{0},\tau_{1},\tau_{2},\tau_{3}].$
Let $\varphi(x)\equiv x\ln x$ with $\varphi(0)\equiv 0$ for $x\in[0;1]$ and let $\tau=(\tau_{0},\tau_{1},\tau_{2},\tau_{3})$. Define
$C_{1}\equiv\left\{\tau:\tau_{i}\geqslant 0,i\in\{0,1,2,3\},\sum\limits^{3}_{i=0}\tau_{i}=1,\tau_{1}+2\tau_{2}+3\tau_{3}=\theta\right\},$
where $\theta\in(0;3)$ and let $f(\tau)\equiv\sum\limits^{3}_{i=0}\tau_{i}\ln\tau_{i}$. According to the Weierstrass theorem [16, p. 134], the function $f(\tau)$ attains its minimum on the compact set $C_{1}$, and we denote this minimum by $m(\theta)$.
Theorem 2.
The set $\Theta_{1}$ is a continuous, everywhere dense, closed set of zero Lebesgue measure for $\theta\neq\frac{3}{2}$ and of full Lebesgue measure for $\theta=\frac{3}{2}$. Its Hausdorff–Besicovitch fractal dimension $\alpha_{0}(\Theta_{1})$ satisfies the inequality
$\alpha_{0}(\Theta_{1})\geqslant-\frac{m(\theta)}{\ln 4}.$
Proof.
Since the set $E[\tau_{0},\tau_{1},\tau_{2},\tau_{3}]$ is continuous and everywhere dense set, the same holds for the set $\Theta_{1}$. The set $\Theta_{1}$ is closed since all its points are limit points. Indeed, for any $\Theta_{1}\ni x_{0}=\Delta^{4}_{a_{1}a_{2}\ldots a_{n}\ldots}$ there exists a sequence $x_{n}=\Delta^{4}_{a_{1}(x_{0})a_{2}(x_{0})\ldots a_{n}(x_{0})}$ such that $\lim\limits_{n\to\infty}x_{n}=x_{0}$.
If $\theta\neq\displaystyle\frac{3}{2}$, then $\Theta_{1}$ contains no normal numbers. Since almost all numbers (in the sense of Lebesgue measure) are normal, then we have $\lambda(\Theta_{1})=0$. On the other hand, for $\theta=\displaystyle\frac{3}{2}$, the following inclusion holds: $\Theta_{1}\supset E\left[\displaystyle\frac{1}{4},\displaystyle\frac{1}{4},\displaystyle\frac{1}{4},\displaystyle\frac{1}{4}\right]$, where $\lambda(E)=1$. Thus, $\lambda(\Theta_{1})=1$.
Let $\tau=(p_{0},p_{1},p_{2},p_{3})$ be such that $f(\tau)=m(\theta)$. Then, according to the Besicovitch–Eggleston formula, we have
$\alpha_{0}(E[p_{0},p_{1},p_{2},p_{3}])=-\frac{\ln p_{0}^{p_{0}}p_{1}^{p_{1}}p_{2}^{p_{2}}p_{3}^{p_{3}}}{\ln 4}=-\frac{f(\tau)}{\ln 4}=-\frac{m(\theta)}{\ln 4}.$
Since $E[p_{0},p_{1},p_{2},p_{3}]\subset\Theta_{1}$, then $\alpha_{0}(\Theta_{1})\geqslant\alpha_{0}(E[p_{0},p_{1},p_{2},p_{3}])=-\displaystyle\frac{m(\theta)}{\ln 4}.$
3. An example of a number from the set $\Theta_{1}$
Let us present an algorithm for constructing a number $x\in E[\tau_{0},\tau_{1},\tau_{2},\tau_{3}]$.
We consider the sequences $\tau_{in}=[\tau_{i}\cdot n]$ and $\tau^{\prime}_{in}=\tau_{i(n+1)}-\tau_{in}$. Clearly,
$\tau_{i(n+1)}]-[\tau_{in}]=\left[[\tau_{in}]+\{\tau_{in}\}+\tau_{i}\right]-[\tau_{in}]=[\tau_{in}]+[\{\tau_{in}\}+\tau_{i}]-[\tau_{in}]=[\{\tau_{in}\}+\tau_{i}]\in\{0,1\}.$
It is evident that
$\frac{\tau_{in}}{n}=\frac{[\tau_{i}\cdot n]}{n}=\frac{\tau_{i}\cdot n-\{\tau_{i}\cdot n\}}{n}=\tau_{i}-\frac{\{\tau_{i}\cdot n\}}{n}\to\tau_{i},\penalty 10000\ n\to\infty.$
We construct the number $x$ as follows. In the first step, we sequentially write $\tau^{\prime}_{01}$ zeros, $\tau^{\prime}_{11}$ ones, $\tau^{\prime}_{21}$ twos, and $\tau^{\prime}_{31}$ threes. After the $k$-th step, in the $(k+1)$-step we append to the alredy written sequence of $\tau^{\prime}_{0k}$ zeros, $\tau^{\prime}_{1k}$ ones, $\tau^{\prime}_{2k}$ twos, and $\tau^{\prime}_{3k}$ threes. As a result, among the first $\sum\limits_{i=0}^{3}\tau_{in}$ symbols of the 44–adic representation of $x$ there are exactly $\tau_{in}$ digits equal to ii. Let $n$ be a sufficiently large natural number. Since for any $x\in\mathbb{R}$ we have $x-1<[x]\leqslant x$, then it follows that
$n=\sum\limits^{3}_{i=0}\tau_{i}\cdot n\geqslant\sum\limits^{3}_{i=0}[\tau_{i}\cdot n]=\sum\limits^{3}_{i=0}\tau_{in},$ $\sum\limits^{3}_{i=0}\tau_{i(n+4)}=\sum\limits^{3}_{i=0}[\tau_{i(n+4)}]\geqslant\sum\limits^{3}_{i=0}\tau_{i}\cdot(n+4)-4=n.$
Then $v^{(n)}_{i}\geqslant\displaystyle\frac{\tau_{in}}{n}\to\tau_{i}$, $v^{(n)}_{i}\leqslant\displaystyle\frac{\tau_{i(n+3)}}{n}=\displaystyle\frac{\tau_{i(n+3)}}{n+3}\cdot\displaystyle\frac{n+3}{n}\to\tau_{i}$ as $n\to\infty$. Hence, $\nu_{i}(x)=\tau_{i}$ for all $i\in\{0,1,2,3\}$.
Let $(s_{k})$ be a sequence of positive numbers such that
$\lim\limits_{k\to\infty}s_{k}=\infty,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \lim\limits_{k\to\infty}\displaystyle\frac{s_{k+1}}{\sum\limits^{k}_{i=1}s_{i}}=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \lim\limits_{k\to\infty}\displaystyle\frac{k}{\sum\limits^{k}_{i=1}s_{i}}=0.$
Let $\|\tau_{in}\|$ be a $(4\times\infty)$ matrix whose elements are the numbers constructed above.
We consider the following form of representation of a real number $x\in[0,1]$:
$\hat{x}=\Delta^{4}_{\underbrace{\underbrace{0\ldots 0}_{[\tau_{01}s_{1}]}\underbrace{1\ldots 1}_{[\tau_{11}s_{1}]}\underbrace{2\ldots 2}_{[\tau_{21}s_{1}]}\underbrace{3\ldots 3}_{[\tau_{31}s_{1}]}}_{\text{1st block}}\ldots\underbrace{\underbrace{0\ldots 0}_{[\tau_{0k}s_{k}]}\underbrace{1\ldots 1}_{[\tau_{1k}s_{k}]}\underbrace{2\ldots 2}_{[\tau_{2k}s_{k}]}\underbrace{3\ldots 3}_{[\tau_{3k}s_{k}]}}_{\text{k-th block}}\ldots}.$ (1)
Theorem 3.
If $\|\tau_{in}\|$ is a $(4\times\infty)$ matrix such that for any natural number $n\in N$ the conditions $\tau_{0n}+\tau_{1n}+\tau_{2n}+\tau_{3n}=1$ and $\tau_{1n}+2\tau_{2n}+3\tau_{3n}=\theta$ are satisfied, then
$\lim\limits_{n\to\infty}r_{n}(\hat{x})=\theta.$
Proof.
Since $\tau_{0k}s_{k}]+[\tau_{1k}s_{k}]+[\tau_{2k}s_{k}]+[\tau_{3k}s_{k}]>\tau_{0k}s_{k-1}+\tau_{1k}s_{k-1}+\tau_{2k}s_{k-1}+\tau_{3k}s_{k-1}=s_{k-4}\to\infty$ as $k\to\infty,$ then the number $\hat{x}$ is constructed correctly.
Let $n$ be a sufficiently large natural number, and suppose the nn‑th digit of the number $\hat{x}$ falls within the kk‑th block. We introduce the following notation:
$A_{k}\equiv\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{k}_{i=1}([\tau_{1i}s_{i}]+2[\tau_{2i}s_{i}]+3[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}},$ $B_{k}\equiv\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}}.$
then as $k\to\infty$ we have
$A_{k}\leqslant\displaystyle\frac{\sum\limits^{k}_{i=1}(\tau_{1i}s_{i}+2\tau_{2i}s_{i}+3\tau_{3i}s_{i})}{\sum\limits^{k}_{i=1}s_{i}}=\displaystyle\frac{\sum\limits^{k}_{i=1}\theta s_{i}}{\sum\limits^{k}_{i=1}s_{i}}=\theta,$
$A_{k}>\displaystyle\frac{\sum\limits^{k}_{i=1}((\tau_{1i}s_{i}-1)+2(\tau_{2i}s_{i}-1)+3(\tau_{3i}s_{i}-1))}{\sum\limits^{k}_{i=1}s_{i}}=\theta-\frac{6k}{\sum\limits^{k}_{i=1}s_{i}}.$
Hence, $A_{k}=\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{k}_{i=1}([\tau_{1i}s_{i}]+2[\tau_{2i}s_{i}]+3[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}}=\theta.$
On the other hand
$B_{k}\leqslant\displaystyle\frac{\sum\limits^{k}_{i=1}(\tau_{0i}s_{i}+\tau_{1i}s_{i}+\tau_{2i}s_{i}+\tau_{3i}s_{i})}{\sum\limits^{k}_{i=1}s_{i}}=\displaystyle\frac{\sum\limits^{k}_{i=1}s_{i}}{\sum\limits^{k}_{i=1}s_{i}}=1,$
$B_{k}>\displaystyle\frac{\sum\limits^{k}_{i=1}(\tau_{0i}s_{i}+\tau_{1i}s_{i}+\tau_{2i}s_{i}+\tau_{3i}s_{i}-4)}{\sum\limits^{k}_{i=1}s_{i}}=1-\frac{4k}{\sum\limits^{k}_{i=1}s_{i}}.$
Therefore $B_{k}=\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}}=1.$
$\frac{\sum\limits^{k+1}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}}=\displaystyle\frac{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}}+$
$\frac{s_{k+1}-(\{\tau_{0(k+1)}s_{k+1}\}+\{\tau_{1(k+1)}s_{k+1}\}+\{\tau_{2(k+1)}s_{k+1}\}+\{\tau_{3(k+1)}s_{k+1}\})}{\sum\limits^{k}_{i=1}s_{i}}\to 1.$
We have
$r_{n}(\hat{x})\geqslant\frac{\sum\limits^{k}_{i=1}([\tau_{1i}s_{i}]+2[\tau_{2i}s_{i}]+3[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}=\frac{A_{k}}{B_{k}}\to\frac{\theta}{1}=\theta.$
Let $B^{\prime}_{k}\equiv\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k+1}_{i=1}s_{i}}.$ Then
$B^{\prime}_{k}=\displaystyle\frac{\sum\limits^{k+1}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k+1}_{i=1}s_{i}}$-
$\frac{s_{k+1}-(\{\tau_{0(k+1)}s_{k+1}\}+\{\tau_{1(k+1)}s_{k+1}\}+\{\tau_{2(k+1)}s_{k+1}\}+\{\tau_{3(k+1)}s_{k+1}\})}{\sum\limits^{k+1}_{i=1}s_{i}}\to 1.$
We have
$r_{n}(\hat{x})\leqslant\frac{\sum\limits^{k+1}_{i=1}([\tau_{1i}s_{i}]+2[\tau_{2i}s_{i}]+3[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}=\frac{A_{k}}{B^{\prime}_{k}}\to\frac{\theta}{1}=\theta.$
Hence, $\lim\limits_{n\to\infty}r_{n}(\hat{x})=\theta.$
Theorem 4.
If $\|\tau_{in}\|$ is a stochastic $(4\times\infty)$ matrix such that for a fixed $j\in{0,1,2,3}$ holds $\lim\limits_{n\to\infty}\tau_{jn}=\lambda_{j},$ then
$\nu_{j}(\hat{x})=\lambda_{j},$
where the number $\hat{x}$ has the form given in (1).
Proof.
We denote $x_{n}=\sum\limits^{n}_{i=1}\tau_{ij}s_{i},$ $y_{n}=\sum\limits^{n}_{i=1}s_{i}.$ Then $\lim\limits_{n\to\infty}y_{n}=\infty,$ $\lim\limits_{n\to\infty}\displaystyle\frac{x_{n+1}-x_{n}}{y_{n+1}-y_{n}}=\lim\limits_{n\to\infty}\displaystyle\frac{\tau_{j(n+1)}s_{n+1}}{s_{n+1}}=\lim\limits_{n\to\infty}{\tau_{j(n+1)}}=\lambda_{j}.$
Therefore, by the Stolz theorem [15, p. 67] we have
$\lim\limits_{n\to\infty}\frac{\sum\limits^{n}_{i=1}\tau_{ji}s_{i}}{\sum\limits^{n}_{i=1}s_{i}}=\lim\limits_{n\to\infty}\frac{x_{n}}{y_{n}}=\lambda_{j}.$
It follows from the proof of the previous theorem that the number $\hat{x}$ is correctly constructed and
$\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{k-1}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}}=1=\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{k+1}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}{\sum\limits^{k}_{i=1}s_{i}}.$
Let $n$ be a sufficiently large natural number, and suppose the nn–th digit of the number $\hat{x}$ falls within the kk–th block. Then
$\frac{\sum\limits^{k}_{i=1}[\tau_{ji}s_{i}]}{\sum\limits^{k}_{i=1}s_{i}}\leqslant\displaystyle\frac{\sum\limits^{k}_{i=1}\tau_{ji}s_{i}}{\sum\limits^{k}_{i=1}s_{i}}\to\lambda_{j}$ and $\frac{\sum\limits^{k}_{i=1}[\tau_{ji}s_{i}]}{\sum\limits^{k}_{i=1}s_{i}}>\displaystyle\frac{\sum\limits^{k}_{i=1}(\tau_{ji}s_{i}-1)}{\sum\limits^{k}_{i=1}s_{i}}=\displaystyle\frac{x_{k}}{y_{k}}-\frac{k}{\sum\limits^{k}_{i=1}s_{i}}\to\lambda_{j},\penalty 10000\ (k\to\infty).$ Hence, $\lim\limits_{k\to\infty}\displaystyle\frac{\sum\limits^{n}_{i=1}[\tau_{ji}s_{i}]}{\sum\limits^{n}_{i=1}s_{i}}=\lambda_{j}.$
$N_{j}(\hat{x},n)\geqslant\frac{\sum\limits^{k}_{i=1}[\tau_{ji}s_{i}]}{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}=\frac{\frac{\sum\limits^{k}_{i=1}[\tau_{ji}s_{i}]}{y_{k}}}{\frac{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+\ldots+[\tau_{3i}s_{i}])}{y_{k}}}\to\frac{\lambda_{j}}{1}=\lambda_{j},\penalty 10000\ k\to\infty,$
$N_{j}(\hat{x},n)\leqslant\frac{\sum\limits^{k+1}_{i=1}[\tau_{ji}s_{i}]}{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+[\tau_{1i}s_{i}]+[\tau_{2i}s_{i}]+[\tau_{3i}s_{i}])}=\frac{\frac{\sum\limits^{k+1}_{i=1}[\tau_{ji}s_{i}]}{y_{k+1}}}{\frac{\sum\limits^{k}_{i=1}([\tau_{0i}s_{i}]+\ldots+[\tau_{3i}s_{i}])}{y_{k+1}}}\to\frac{\lambda_{j}}{1}=\lambda_{j},\penalty 10000\ k\to\infty.$
Therefore, $\nu_{j}(\hat{x})=\lambda_{j}$. ∎
Theorem 5.
Let $(s^{(r)}_{k})$ for $r\in\{1,2\}$ be sequences of positive numbers such that $\lim\limits_{k\to\infty}s^{(r)}_{k}$$\infty$. Let $\|p^{(1)}\|=\|p^{(1)}_{in}\|$, $\|p^{(2)}\|=\|p^{(2)}_{in}\|$ be stochastic $(4\times\infty)$ matrices. Let
$x(\|p^{(r)}\|;\|s^{(j)}_{k}\|)=\Delta^{4}_{\underbrace{\underbrace{0\ldots 0}_{[p^{(r)}_{01}s^{(j)}_{1}]}\underbrace{1\ldots 1}_{[p^{(r)}_{11}s^{(j)}_{1}]}\underbrace{2\ldots 2}_{[p^{(r)}_{21}s^{(j)}_{1}]}\underbrace{3\ldots 3}_{[p^{(r)}_{31}s^{(j)}_{1}]}}_{\text{1st block}}\ldots\underbrace{\underbrace{0\ldots 0}_{[p^{(r)}_{0k}s^{(j)}_{k}]}\underbrace{1\ldots 1}_{[p^{(r)}_{1k}s^{(j)}_{k}]}\underbrace{2\ldots 2}_{[p^{(r)}_{2k}s^{(j)}_{k}]}\underbrace{3\ldots 3}_{[p^{(r)}_{3k}s^{(j)}_{k}]}}_{\text{k-th block}}\ldots}.$
[…]
If $\overline{\lim\limits_{n\to\infty}}\sum\limits^{3}_{i=0}|p^{(1)}_{in}-p^{(2)}_{in}|>0$, then $x(\|p^{(1)}\|;\|s^{(1)}_{k}\|)\neq x(\|p^{(2)}\|;\|s^{(2)}_{k}\|)$.
Proof.
[…]
Now let $\overline{\lim\limits_{n\to\infty}}\sum\limits^{3}_{i=0}|p^{(1)}_{in}-p^{(2)}_{in}|>0$ and $x(\|p^{(1)}\|;\|s^{(1)}_{k}\|)=x(\|p^{(2)}\|;\|s^{(2)}_{k}\|)$. Then all nn‑th blocks of the numbers $x(\|p^{(1)}\|;\|s^{(1)}_{k}\|)$ and $x(\|p^{(2)}\|;\|s^{(2)}_{k}\|)$ are equal, that is, $p^{(1)}_{in}s^{(1)}_{n}]=[p^{(2)}_{in}s^{(2)}_{n}$], $n\in N$, $i\in\{0,1,2,3\}$. Hence, $|p^{(1)}_{in}-p^{(2)}_{in}|s^{(1)}_{n}<1$ for all $n\in N$, $i\in\{0,1,2,3\}$, which is only possible if $\lim\limits_{n\to\infty}|p^{(1)}_{in}-p^{(2)}_{in}|=0$. This leads to a contradiction. ∎